A bob of mass M is suspended by a masslessstring of length L . The horizontal velocityat position 4 is just sufficient to make itreach the point The angle θ at which thespeed of the bob is half of that at A, satisfiesa)theta=pi/4 b)pi/4

Grade:11

1 Answers

Arun
6 years ago
If you see in the diagram,
You get
(1/2)* 5 mgl = (1/2)* m* (5gl/4) + mgl (1- cosθ)
On solving
Cos θ = - 7/8
Hence
θ > 3π4
θπ
Hope it helps